### v1

@@ 3. $\bullet$
      For any pair $(s,i)$, either $\tau^{*}\neq\tau_{si}$ (Type-1 forgery) where $s\in[l]$ and $i\in[q_{s}]$, or there exists a pair $(s,i)$ such that $\tau^{*}=\tau_{si}$ (Type-2 forgery) but $\mathbf{m}^{*}\notin V_{si}$.
@@ 1. $\bullet$
    KeyGen ($1^{n}$, pp): Let $q=\text{poly}(n)$, $k=\text{poly}(n)$ be an even integer, with $q\geq(nk)^{3}$, $h=\left\lfloor\frac{n}{6\log q}\right\rfloor$, and $V=\sqrt{2nk\log q}\cdot\log n$. Sample $\bm{\alpha}_{1},...,\bm{\alpha}_{k}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$, and let the public parameters be
    $\textsf{pp}=\{q,k,h,V,\bm{\alpha}_{1},...,\bm{\alpha}_{k}\}.$
    Furthermore, the tag space is $\mathcal{T}=\{0,1\}^{k}$, the message space $\mathcal{M}=\mathbb{F}_{2}^{k}$, and the signature space $\Sigma=\mathbb{Z}^{\ell n}$, where $\ell$ is the number of users in ring $R$ and $k$ bounds the maximum homomorphic operations.
@@
      Set the public key as $pk_{s}=A_{s}$ and the secret key as $sk_{s}=T_{s}$, then output the key pair $(pk_{s},sk_{s})=(A_{s},T_{s})$ for user $s$.
@@ 2. $\bullet$
    HR-Sign$(pk_{s},sk_{s},R_{\ell},\tau,\mathbf{m})$: Given the public-private key pair of user $s$, $(pk_{s},sk_{s})=(A_{s},T_{s})$, and a ring $R_{\ell}$ consisting of $l$ users’ public keys, we simplify notation by letting $R_{\ell}=\{A_{1},\ldots,A_{\ell}\}$ where $pk_{s}\in R_{\ell}$. For a tag $\tau=(\tau_{1},...,\tau_{k})\in\{0,1\}^{k}$ and message $\mathbf{m}=(m_{1},...,m_{k})\in\mathbb{F}_{2}^{k}$, the signature is generated as follows:
@@
      If $\mathbf{u}$ and $\mathbf{v}$ contain nonzero vectors (note: by Theorem 1, at most one of $\mathbf{u}$ or $\mathbf{v}$ is nonzero), set $\mathbf{t}=\sum_{j=1}^{k}(-1)^{\tau_{j}}m_{j}\bm{\alpha}_{j}$ and compute
@@ 4. $\bullet$
    HR-Verify$(R_{\ell},\tau,\mathbf{m},\sigma)$: On input the ring $R_{\ell}$, the tag $\tau$, the message $\mathbf{m}$ and the signature $\sigma=(\mathbf{e},L_{R_{\ell}})$,where $L_{R_{\ell}}=(i_{1},i_{2},...i_{\ell})$, the verification process is as follows:
@@
        $\parallel\mathbf{e}\parallel\leq 2kV\sqrt{2kln}$;
@@ 4.2 Correctness
    If $\sigma=(\mathbf{e},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(pk_{s},sk_{s},R_{\ell},\tau,\mathbf{m})$. Next, we show that the verification conditions will hold with overwhelming probability.
@@
    implying $V\geq\parallel\widetilde{T}_{s}\parallel\cdot\omega(\sqrt{\log n})$. Thus, Lemma 5 ensures that, with overwhelming probability, $\parallel\mathbf{e}\parallel=\parallel\mathbf{e}(\mathbf{u})+\mathbf{e}(\mathbf{v})\parallel\leq 2V\sqrt{\ell n}<2kV\sqrt{2k\ell n}$. Moreover, Lemma 2 yields the congruence relation:
@@
    $\sigma_{j}=(\mathbf{e}_{j},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(pk_{s},sk_{s},R_{\ell},\tau,\mathbf{m}_{j}).$
@@
[…]
@@
    $\Big{\|}\sum_{j=1}^{p}c_{j}\mathbf{e}_{j}\Big{\|}\leq 2pV\sqrt{\ell n}\leq 2kV\sqrt{\ell n}\leq 2kV\sqrt{2k\ell n}.$
@@
[…]
@@ 5.1 Anonymity
    -Sign query: The adversary $\mathcal{A}_{1}$ selects and sends to $\mathcal{B}_{1}$ the tuple $(s,R_{\ell},V_{si})$, where the message subspace $V_{si}$ has basis vectors $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k}$ and satisfies $pk_{s}\in R_{\ell}$, with $1\leq\ell\leq q_{E}$, $1\leq i\leq q_{s}$, and $q_{s}$ denoting the maximum number of queries that user $s$ can make on message subspace $V_{si}$ under ring $R_{\ell}$. $\mathcal{B}_{1}$ first generates a random tag $\tau_{si}=(\tau_{si}^{(1)},...,\tau_{si}^{(k)})\stackrel{{\scriptstyle\$}}{{\leftarrow}}\{0,1\}^{k}$ for $V_{si}$ and produces the signature as follows:
@@
    The challenger $\mathcal{B}_{1}$ returns the tuple $(s,R_{\ell},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to the adversary $\mathcal{A}_{1}$, where $pk_{s}\in R_{\ell}$, $l\in[q_{E}]$, $1\leq i\leq q_{s}$, and $1\leq j\leq k$.
@@
    where $s_{0}$ and $s_{1}$ are indices satisfying $pk_{s_{0}}\in R_{\ell^{*}}$ and $pk_{s_{1}}\in R_{\ell^{*}}$, and $(\tau^{*},\mathbf{m}^{*})$ is the data message to be signed by the ring $R_{\ell^{*}}$. $\mathcal{B}_{1}$ randomly selects a bit $b\in\{0,1\}$, generates a signature $\mathbf{e}_{b}^{*}$ following the same procedure as the signing query, and provides $\mathbf{e}_{b}^{*}$ to $\mathcal{A}_{1}$. Finally, $\mathcal{A}_{1}$ outputs a bit $b^{\prime}$.
@@ 5.2 Unforgeability
Assuming the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem is hard, where $\ell$ denotes the size of the challenge ring and $\beta=4kV\sqrt{2k\ell n}$, then the $\mathcal{LHRS}$ scheme constructed in Section 4 achieves unforgeability against against insider corruption.
More specifically, for a polynomial-time adversary $\mathcal{A}_{2}$ who is allowed to access rings of maximum size $q_{E}$ and can make at most $q_{s}$ signing queries per user $s$ in any ring (relative to the message subspace), we can construct a polynomial-time algorithm $\mathcal{B}_{2}$ that solves the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem with the following advantage:
$\mathbf{Adv}_{q,h,\ell n,2V}^{\mathrm{SIS}}(\mathcal{B}_{2})\geq\frac{\mathbf{Adv}_{\mathrm{LHRS}}^{\mathrm{UNF}}(\mathcal{A}_{2})}{2q_{E}\binom{q_{E}}{q_{E}/2}}.$
@@ 1. $\bullet$
    Setup: In the system, the parameters $q$, $k$, $h$, and the size of $V$ follow the same configuration as specified in Section 4. Sample $k$ vectors: $\mathbf{x}_{1},\dots,\mathbf{x}_{k}\xleftarrow{\$}\textsf{SampleDom}(1^{\ell n},\frac{1}{\sqrt{k/2}}V)$. Let $\bm{\alpha}_{1}=A_{R_{\mathbf{w}}}\mathbf{x}_{1}$, $\dots$, $\bm{\alpha}_{k}=A_{R_{\mathbf{w}}}\mathbf{x}_{k}$. Observe that $\frac{1}{\sqrt{k/2}}V\geq\omega\big{(}\sqrt{\log(q_{E}n)}\big{)}\geq\omega\big{(}\sqrt{\log(\ell n)}\big{)}$. According to Lemma 3, the vectors $\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}$ are statistically close to the uniform distribution over $\mathbb{Z}_{q}^{h}$. $\mathcal{B}_{2}$ then sends the public parameters $\textsf{pp}=\{q,k,h,V,\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\}$ to $\mathcal{A}_{2}$. Finally, $\mathcal{B}_{2}$ generates public-private key pairs for all $q_{E}$ users in the following manner:
@@ 2. $\bullet$
$\left\|\mathbf{e}_{si}^{(j)}\right\|=\left\|\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})\right\|\leq 2V\sqrt{2k\ell^{*}n}\leq 2kV\sqrt{2k\ell^{*}n}.$
@@
$\left\|\mathbf{x}\right\|=\left\|\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\mathbf{x}_{i}\right\|\leq k\cdot\frac{V}{\sqrt{k/2}}\sqrt{\ell n}=V\sqrt{2k\ell n}\leq 2kV\sqrt{2k\ell n},$
@@
$\left\|\mathbf{e}^{*}-\mathbf{x}\right\|\leq 4kV\sqrt{2k\ell n}.$
@@
Worst-case to Average-case Reduction. Building upon Theorem 2 in [38], the computational complexity of the $\mathrm{SIS}_{q,n,m,\beta}$ can be shown to be polynomially equivalent to worst-case approximations of the Shortest Independent Vectors Problem (SIVP), given that the modulus satisfies $q\geq\beta\cdot\omega(\sqrt{n\log n})$. Under this condition, the approximation factor for SIVP is on the order of $\beta\cdot\tilde{O}(\sqrt{n})$. In our construction, we require that $q\geq(nk)^{3}$. Therefore, there exists a sufficiently large $q$ that meets the conditions of Theorem 2, thereby ensuring its applicability.
6 Conclusions and Open Problems
@@
This scheme still presents several promising directions for further research and improvement: (1) First, exploring the construction of novel schemes with higher computational efficiency and more compact signature sizes; (2) Investigating the feasibility of designing homomorphic ring signature schemes that support polynomial or fully homomorphic computations; (3) Examining whether homomorphic ring signature schemes with tight or almost-tight security can be constructed under standard lattice assumptions.

### v3

@@ 3. $\bullet$
      For any pair $(s,i)$, either $\tau^{*}\neq\tau_{si}$ (Type-1 forgery) where $s\in[\ell]$ and $i\in[q_{s}]$, or there exists a pair $(s,i)$ such that $\tau^{*}=\tau_{si}$ (Type-2 forgery) but $\mathbf{m}^{*}\notin V_{si}$.
@@ 1. $\bullet$
    Setup ($1^{n}$, pp): Let $q=\text{poly}(n)$, $k_{0}=\text{poly}(n)$ (bounding maximum homomorphic operations), and $k=\text{poly}(n)$ be such that $k$ is even and $k_{0}<k$, with $q\geq(nkk_{0})^{3}$. Define $h=\left\lfloor\frac{n}{6\log q}\right\rfloor$ and $V=\sqrt{2nk\log q}\cdot\log n$. Sample $\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$, and set the public parameters as $\textsf{pp}=\{q,k_{0},k,h,V,\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\}$. The tag space is $\mathcal{T}=\{0,1\}^{k}$, the message space $\mathcal{M}=\mathbb{F}_{2}^{k}$, and the signature space $\Sigma=\mathbb{Z}^{\ell n}$ where $\ell$ denotes the number of users in ring $R$.
@@
      Set the public key as $\textsf{pk}_{s}=A_{s}$ and the secret key as $\textsf{sk}_{s}=T_{s}$, then output the key pair $(\textsf{pk}_{s},\textsf{sk}_{s})=(A_{s},T_{s})$ for user $s$.
@@ 2. $\bullet$
[…]
@@
      If $\mathbf{u}$ and $\mathbf{v}$ contain zero vectors (note: by Theorem 1, at most one of $\mathbf{u}$ or $\mathbf{v}$ is zero), set $\mathbf{t}=\sum_{j=1}^{k}(-1)^{\tau_{j}}m_{j}\bm{\alpha}_{j}$ and compute
@@ 4. $\bullet$
[…]
@@
        $\parallel\mathbf{e}\parallel\leq 2k_{0}V\sqrt{2k\ell n}$;
@@ 4.2 Correctness
    If $\sigma=(\mathbf{e},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R_{\ell},\tau,\mathbf{m})$. Next, we show that the verification conditions will hold with overwhelming probability.
@@
    implying $V\geq\parallel\widetilde{T}_{s}\parallel\cdot\omega(\sqrt{\log n})$. Thus, Lemma 5 ensures that, with overwhelming probability, $\parallel\mathbf{e}\parallel=\parallel\mathbf{e}(\mathbf{u})+\mathbf{e}(\mathbf{v})\parallel\leq 2V\sqrt{\ell n}<2k_{0}V\sqrt{2k\ell n}$. Moreover, Lemma 2 yields the congruence relation:
@@
    $\sigma_{j}=(\mathbf{e}_{j},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R_{\ell},\tau,\mathbf{m}_{j}).$
@@
    $\textsf{HR-Verify}(R_{\ell},\tau,\sum_{j=1}^{p}c_{j}\mathbf{m}_{j},\textsf{Combine}(R_{\ell},\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{p}))\rightarrow 1.$
@@
    $\Big\|\sum_{j=1}^{p}c_{j}\mathbf{e}_{j}\Big\|\leq 2pV\sqrt{\ell n}\leq 2k_{0}V\sqrt{\ell n}\leq 2k_{0}V\sqrt{2k\ell n}.$
@@
    $\textsf{HR-Verify}(R_{\ell},\tau,\sum_{j=1}^{p}c_{j}\mathbf{m}_{j},\textsf{Combine}(R_{\ell},\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{p}))\rightarrow 1.$
@@ 5.1 Anonymity
    -Sign query: The adversary $\mathcal{A}_{1}$ selects and sends to $\mathcal{B}_{1}$ the tuple $(s,R_{\ell},V_{si})$, where the message subspace $V_{si}$ has basis vectors $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k_{0}}$ and satisfies $\textsf{pk}_{s}\in R_{\ell}$, with $1\leq\ell\leq q_{E}$, $1\leq i\leq q_{s}$, and $q_{s}$ denoting the maximum number of queries that user $s$ can make on message subspace $V_{si}$ under ring $R_{\ell}$. $\mathcal{B}_{1}$ first generates a random tag $\tau_{si}=(\tau_{si}^{(1)},...,\tau_{si}^{(k)})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{k}$ for $V_{si}$ and produces the signature as follows:
@@
    The challenger $\mathcal{B}_{1}$ returns the tuple $(s,R_{\ell},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to the adversary $\mathcal{A}_{1}$, where $\textsf{pk}_{s}\in R_{\ell}$, $\ell\in[q_{E}]$, $1\leq i\leq q_{s}$, and $1\leq j\leq k_{0}$.
@@
    where $s_{0}$ and $s_{1}$ are indices satisfying $\textsf{pk}_{s_{0}}\in R_{\ell^{*}}$ and $\textsf{pk}_{s_{1}}\in R_{\ell^{*}}$, and $(\tau^{*},\mathbf{m}^{*})$ is the data message to be signed by the ring $R_{\ell^{*}}$. $\mathcal{B}_{1}$ randomly selects a bit $b\in\{0,1\}$, generates a signature $\mathbf{e}_{b}^{*}$ following the same procedure as the signing query, and provides $\mathbf{e}_{b}^{*}$ to $\mathcal{A}_{1}$. Finally, $\mathcal{A}_{1}$ outputs a bit $b^{\prime}$.
@@ 5.2 Unforgeability
Next, we demonstrate that in schemes with smaller ring sizes (for example, less than or equal to $\mathcal{O}(\log n)$), the aforementioned $\mathcal{LHRS}$ satisfies unforgeability.
@@
Assuming the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem is hard, where $\ell$ denotes the size of the challenge ring and $\beta=4k_{0}V\sqrt{2k\ell n}$, then the $\mathcal{LHRS}$ scheme constructed in Section 4 achieves unforgeability against against insider corruption.
More specifically, for a polynomial-time adversary $\mathcal{A}_{2}$ who is allowed to access rings of maximum size $q_{E}(\leq\mathcal{O}(\log n))$ and can make at most $q_{s}$ signing queries per user $s$ in any ring (relative to the message subspace), we can construct a polynomial-time algorithm $\mathcal{B}_{2}$ that solves the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem with the following advantage:
@@ 1. $\bullet$
    Setup: In the system, the parameters $q$, $k$, $h$, and the size of $V$ follow the same configuration as specified in Section 4. Sample $k$ vectors: $\mathbf{x}_{1},\dots,\mathbf{x}_{k}\xleftarrow{\mathdollar}\textsf{SampleDom}(1^{\ell n},\frac{1}{\sqrt{k/2}}V)$. Let $\bm{\alpha}_{1}=A_{R_{\mathbf{w}}}\mathbf{x}_{1}$, $\dots$, $\bm{\alpha}_{k}=A_{R_{\mathbf{w}}}\mathbf{x}_{k}$. Observe that $\frac{1}{\sqrt{k/2}}V\geq\omega\big(\sqrt{\log(q_{E}n)}\big)\geq\omega\big(\sqrt{\log(\ell n)}\big)$. According to Lemma 3, the vectors $\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}$ are statistically close to the uniform distribution over $\mathbb{Z}_{q}^{h}$. $\mathcal{B}_{2}$ then sends the public parameters $\textsf{pp}=\{q,k_{0},k,h,V,\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\}$ to $\mathcal{A}_{2}$. Finally, $\mathcal{B}_{2}$ generates public-private key pairs for all $q_{E}$ users in the following manner:
@@ 2. $\bullet$
$\left\|\mathbf{e}_{si}^{(j)}\right\|=\left\|\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})\right\|\leq 2V\sqrt{2k\ell^{*}n}\leq 2k_{0}V\sqrt{2k\ell^{*}n}.$
@@
$\left\|\mathbf{x}\right\|=\left\|\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\mathbf{x}_{i}\right\|\leq k\cdot\frac{V}{\sqrt{k/2}}\sqrt{\ell n}=V\sqrt{2k\ell n}\leq 2k_{0}V\sqrt{2k\ell n},$
@@
$\left\|\mathbf{e}^{*}-\mathbf{x}\right\|\leq 4k_{0}V\sqrt{2k\ell n}.$
@@
Analysis of the Security-Loss Factor. The assumption $q_{E}\leq\mathcal{O}(\log n)$ implies that the binomial coefficient $\binom{q_{E}}{q_{E}/2}$ is at most polynomial in $n$. Specifically, using the approximation $\binom{2k}{k}\sim 4^{k}/\sqrt{\pi k}$, the entire denominator $2q_{E}\cdot\binom{q_{E}}{q_{E}/2}$ is of order $\mathcal{O}(\mathsf{poly}(n))$. Therefore, the reduction only incurs a polynomial loss in the advantage. A non-negligible success probability for the forger $\mathcal{A}_{2}$ thus translates to a non-negligible advantage for the SIS solver $\mathcal{B}_{2}$.
Worst-case to Average-case Reduction. Building upon Theorem 2 in [38], the computational complexity of the $\mathrm{SIS}_{q,n,m,\beta}$ can be shown to be polynomially equivalent to worst-case approximations of the Shortest Independent Vectors Problem (SIVP), given that the modulus satisfies $q\geq\beta\cdot\omega(\sqrt{n\log n})$. Under this condition, the approximation factor for SIVP is on the order of $\beta\cdot\tilde{O}(\sqrt{n})$. In our construction, we require that $q\geq(nkk_{0})^{3}$. Therefore, there exists a sufficiently large $q$ that meets the conditions of Theorem 2, thereby ensuring its applicability.
6 Conclusions
@@
This scheme still presents several promising directions for further research and improvement: (1) First, explore the construction of novel schemes featuring short signatures (where the signature size is independent of the ring size) with reasonable computational efficiency; (2) Investigating the feasibility of designing homomorphic ring signature schemes that support polynomial or fully homomorphic computations; (3) Examining whether homomorphic ring signature schemes with tight or almost-tight security can be constructed under standard lattice assumptions.
Acknowledgments
We gratefully acknowledge the support and encouragement of this work from the Academy of Interdisciplinary Studies, Renmin University of China, and Great Bay University.
Author contributions
These authors contributed equally to this work.
Funding
This work was supported by Information Security School-Enterprise Joint Laboratory(Dongguan Institute for Advanced Study, Great Bay University, NO.H24120002).
Availability of data and materials
No data were used in the present study. Thus, there is no relevant data to share or report regarding availability.